# Low-hallucination Synthetic Captions for Large-Scale Vision-Language Model Pre-training

### v1

@@ 3.2 Knowledge-enriching SFT
Previous recaption models (e.g.(Li et al., 2023a; Yu et al., 2024)) often produce generalized and imprecise descriptions, such as "a man performing on stage" in Figure 1. This type of descriptive information has limited informational value and offers little assistance in training vision language models. Hence, in our work, we place significant emphasis on the integration of valuable information. We believe that valuable information is mainly reflected in two aspects: one is the richness of descriptive details and the other is the knowledge content of the description. The former is easy to understand, while the latter is manifested in the supplementation of external knowledge related to the image. For example, in Figure 1, in the case of "oxalic acid", we hope to describe it as follows: The image shows the molecular chemical structure of oxalic acid, with the chemical formula $C_{2}H_{2}O_{4}$. The oxalic acid molecule consists of two carbon atoms, two hydrogen atoms, and four oxygen atoms. In the structural formula, the two carbon atoms are connected by single bonds, and each carbon atom is attached to a hydroxyl group ($OH$) and a carbonyl group ($C=O$), which are the typical structural features of oxalic acid.
Figure 2: A GPT-4o prompt designed to construct image captions rich in descriptive details. In the prompt, the red-highlighted portion lays stress on the methodology of knowledge injection. Furthermore, it guarantees that the ultimately generated caption incorporates a rich body of world knowledge via a two-step procedure.
By constructing SFT data, captions rich in descriptive details can be generated, enabling the model to learn to produce longer captions that contain as much specific information as possible through fine-tuning. Therefore, we carefully constructed a complex prompt and obtained training data using GPT-4o, see Section 3.3. Note that although GPT-4o can understand the requirements of the prompt, the output data still contain a substantial amount of hallucinations. As a result, we perform a manual review. Moreover, to incorporate knowledge into captions more effectively, we screened the original image-text pairs. In addition to general alt-text, we selected object-related data with a higher informational value [^1] In our study, we primarily regulate this aspect through the meticulous selection of alt-text data sources. For example, we deliberately choose a diverse range of encyclopedia data. This allows the descriptions in the captions to focus more on specific object descriptions, such as the "tragocephala pretiosa" in Figure 1. After reinforcement of the above two aspects, we construct knowledge-enhanced supervised fine-tuning data capable of supporting the training of image captions with rich content.
@@ 3.3 Training Pipeline
As mentioned above, our comprehensive training pipeline consists of two sequential stages: knowledge-enriching SFT followed by continuous DPO. In this section, we introduce these stages specifically. The general pipeline of our work is shown in Figure 3. We adopt Qwen2-VL-7B (Wang et al., 2024b) as the base model and use the LLaMA Factory (Zheng et al., 2024) platform to carry out the overall training process.
Figure 3: The illustration of our recaptioning pipeline
Knowledge-enriching SFT. We sampled raw image-text pairs from open-source multimodal datasets including CC3M (Sharma et al., 2018), CC12M (Changpinyo et al., 2021), Datacamp (Gadre et al., 2023) Wukong (Gu et al., 2022), and Wikipeda (Srinivasan et al., 2021). Using the prompt shown in Figure 2, we employed GPT-4o to extract the image-alttext-caption data. Following a meticulous manual review process, $43,408$ data entries were retained. Subsequently, we performed LoRA fine-tuning on the Qwen2-VL-7B model. Specifically, the learning rate was set at $51.0e-$ 5, the global batch size was configured as $128$, and the number of training epochs was established at $10$. Through this process, we eventually acquired our Recaption-SFT model.
Continuous DPO. As outlined in Section 3.1, our methodology begins with training a DPO intermediate model. We first randomly sample $300,000$ image-text pairs and then employ the Recaption-SFT model for inference. To induce controlled stochasticity, we systematically modulate three sampling hyperparameters, TOPP=$1.0$, TOPK=$20$, and TEMPERATURE=$1.0$. Each image-text pair undergoes eight parallel inference iterations, thereby generating eight divergent candidates. We construct preference-ordered pairs by selecting the best and worst outputs from each set with a critic model  [^2] It is an in-house expert system that primarily makes point-by-point judgments by dissecting the details in captions and then assesses the hallucination level of the details. The main pipeline is similar to that of Section 4.1, but it does not use GPT-4o.
To eliminate length-induced training bias, we implement length-balanced sampling that retains $218k$ high-quality pairs, enforcing alignment of length distribution between positive and negative samples. Ultimately, the curated dataset was used for DPO training on the Recaption-SFT model, resulting in the DPO intermediate model.
Upon acquiring the intermediate model, we further sampled $200k$ data instances from the data sources and used this intermediate model to generate preference data with the same hyperparameters and the judge pipeline. Following the length-balancing strategy, we ultimately retained $139k$ data instances for continuous DPO training. All hyperparameters at different stages are listed in Table 2.
@@ 4 Experiments
This section outlines our quality assessment criteria for recaption data and elaborates on our work from two aspects: the intrinsic quality of the dataset and its effectiveness in large-scale vision language model pre-training.
@@
Figure 4: Performance of different DPO strategies. The horizontal axis represents the scale of the training data, and the vertical axis represents the proportion of hallucination-free captions on the validation set (in percentage). The CDPO strategy uses the same training data as DPO in the early stage of training and resamples $139k$ new data after training $218k$ data.
@@ 4.1 Evaluation Metrics
We believe that the quality measurement of recaption data encompasses two aspects. One is the assessment of harmful information, namely hallucinations. This information should be minimized as much as possible. The other aspect is the measurement of content richness, which can usually be verified by the effectiveness of model training. Therefore, in terms of the quality of the recaption data itself, we pay more attention to the indicators of hallucinations, while the measurement of its richness will be reflected through the training effect of large-scale vision language models in Section 4.3. We take advantage of the hallucination assessment method of CIEM (Hu et al., 2023) and leverage the existing large language model, GPT-4o, to decompose potential detailed elements within the caption. Subsequently, we use GPT-4o to make judgments about the presence of hallucinations. Finally, we count the number of hallucinations in a caption. We define the proportion of captions that do not contain hallucinatory details as the non-hallucination rate, and the proportion of captions that contain at most two hallucinatory details as the low-hallucination rate.
@@ 4.2 Performance of Recaption Model
We first report the evaluation results of our recaption model on open-source data, which come mainly from Wukong (Gu et al., 2022). As illustrated in Figure 4, the base model (Qwen2-VL-7B) exhibits a considerably severe level of hallucination. Specifically, its non-hallucination rate ($48.33\%$) and low-hallucination rate ($87.87\%$) are quite low. Following Supervised Fine-Tuning, the low-hallucination rate increases significantly. Subsequently, with the application of DPO, both the non-hallucination rate and the low-hallucination rate are substantially improved, by $34.03\%$ and $10.21\%$, respectively. Furthermore, we observe that DPO demonstrates a notable scaling law during the initial stage of the recaption task. However, the performance plateaus once the data size attains a certain scale. Our CDPO algorithm can further enhance both the non-hallucination rate (by $+3.4\%$) and the low-hallucination rate (by $+0.5\%$).
Meanwhile, we also use our model to infer on the same image-text pairs as other recaption works, e.g. Capfusion (Yu et al., 2024) and Recap-DataComp-1B (Li et al., 2024b), and report the results in Table 1. For the same original alt-text data, the recaption data generated by our model contain significantly more details than that of Capfusion and have much fewer hallucinations than Recap-DataComp-1B. Specifically, when dealing with the same data, compared to Capfusion, the number of details in the captions of our model increases by $267\%$, the average length of our captions is $3.5$ times longer than that of Capfusion, and the rate of hallucinatory details decreases by $8.6\%$. Moreover, compared to Recap-DataComp-1B, the non-hallucination rate increases by $43.6\%$. Note that since the length of captions in Capfusion is short, its non-hallucination rate is quite high. However, because of the lack of details, its value is limited.
@@
Table 3: Experimental results on 15 vision-language tasks. Avg. is the average score of all vision-language tasks. The None-Hallu. Rate is a hallucination assessment metric introduced in Section 4.1. We randomly sampled 1000 pieces of DataCamp data to evaluate the level of hallucination. The text in bold represents the optimal performance for the task under the 20M setting, whereas the text with an underline signifies further enhancement when the entire dataset is utilized. The $\ast$ indicates that a unified lite supervised fine-tuning (SFT) (Wang et al., 2023c; Li et al., 2024a) has been performed on the basis of pre-training. Note that our main purpose is to verify the impact of different caption datasets on the training of VLM models. Therefore, the above results are comparative experiments conducted under a fixed data setting (20M) without carefully supervised fine-tuning and reinforcement learning and are not comparable to SOTA models.
@@ 4.3 Performance of Recaption Data
Vision-Language Model Pre-Training. We first validate the effectiveness of our data set in the pre-training of vision language models. The aim is to demonstrate that, compared to alt-text and other recaption data available on the market, Hunyuan-Recap100M can better assist in the training of vision language models and achieve better performance in both general open source multimodal tasks and in-house knowledge-related test sets. In the VLM framework, we select the LLaVA (Liu et al., 2023) architecture, with SigLIP (Zhai et al., 2023a) as visual encoder and Hunyuan-7B as the base large language model. To make a fair comparison of the impact of caption data, we sample $20M$ data points from alt-text/recaption data. Note that the alt-text group experiment uses the same images as Hunyuan-Recap100M, while Recap-Datacamp-1B uses randomly sampled data [^3] Since our data has undergone knowledge screening, its distribution is not exactly the same as that of DataCamp. As shown in Table 3, on 15 open source multimodal tasks, the VLM model trained with Hunyuan-Recap100M significantly outperforms those trained with alt-text or Recap-Datacamp-1B. Especially in hallucination-related tasks, Hunyuan-Recap100M shows a substantial lead over Recap-Datacamp-1B. Additionally, we evaluated the model’s convergence behavior through loss function monitoring. When using identical supervised fine-tuning datasets, models pre-trained with Hunyuan-Recap100M demonstrated markedly improved convergence properties.
In addition, to further measure the improvement in the perceptive level of the VLM caused by Hunyuan-Recap100M, we manually collected pictures with different visual characteristics from 20 domains, including animals, plants, clothing, landmarks, etc., totaling 1,197 pieces of data with manually annotated labels. We conducted an evaluation of the perceptive ability of the VLM in this set of data and the results are shown in Table 4. The VLM model trained with Hunyuan-Recap100M significantly outperforms other counterparts in 18 of the 20 categories. In general, it brings at least a $6.8\%$ improvement on average. After fully applying Hunyuan-Recap100M, the model has shown significant improvements in 17 of 20 perception tasks, with an additional average improvement of $9\%$.
@@
Alt-text(20M) 34.1 35.9 23.4 22.4 26.2 22.7 24.5 34.5 23.3 27.1 25.5 27.5 54.2 27.3 22.2 65.5 32.6 29.4 20 27.3 30.0
Recap-DataComp-1B(20M) 13.6 19.7 18.9 4.5 29.9 23.2 21.6 33.9 12.2 10.0 12.7 17.0 49.2 13.6 13.1 57.9 19.6 19.6 30.4 17.6 21.8
Hunyuan-Recap100M(20M) 25.0 38.1 26.3 28.4 35.4 41.2 36.5 62.5 25.0 18.5 33.3 40.4 60.7 31.8 31.7 72.4 27.3 34.0 36.5 31.4 36.8
Hunyuan-Recap100M 40.0 44.3 39.0 34.4 53.8 38.1 51.9 63.6 27.5 35.1 31.5 51.0 66.1 50.7 50.8 79.3 46.3 39.6 36.4 38.2 45.8
Table 4: Experimental results on identification of 20 common vision object categories. The text in bold represents the optimal performance for the task under the 20M setting, whereas the text with an underline signifies further enhancement when the entire dataset is utilized.
Text-to-Image Model Training We also verify the effectiveness of Hunyuan-Recap100M on the text-to-image generation task. We perform fine-tuning on the technical stack of Hunyuan-DiT (Li et al., 2024d) in order to demonstrate that our data can help generate specific visual images. As evidenced in Table 5, the publicly released Hunyuan-DiT model can significantly achieve a lower FID and a higher CLIP score on the MSCOCO dataset, as well as our internal perceptual test, when using our recaption data. Furthermore, after fine-tuning the data with Hunyuan-Recap100M, there is a further improvement in both the FID and the CLIP Score. This finding indicates that our data can offer considerable assistance in the realm of world perception and also enable text-to-image models to achieve better performance in terms of authenticity. We present realistic cases in Section B.
@@
Table 5: Text-to-Image Evaluation of Hunyuan-DiT on our internal perceptual test and MSCOCO-30K, w/wo fine-tuned with our recaption data. Note that, due to computational resource limitations, we only sampled $2M$ data from Hunyuan-Recap100M for LoRA fine-tuning.

### v2

@@ 3.2 Knowledge-enriching SFT
Previous recaption models (e.g. (Li et al., 2023a; Yu et al., 2024)) often produce generalized descriptions lacking specificity, such as "a man performing on stage" in Figure 1. This type of descriptive information has limited informational value, thus offering little assistance in training vision language models. Hence, our work significantly emphasizes the integration of valuable information. We define valuable information as comprising two main aspects: the richness of descriptive details and the knowledge content of the description. The latter involves the incorporation of external knowledge related to the image. For example, considering "oxalic acid" in Figure 1, we aim for descriptions such as: The image shows the molecular chemical structure of oxalic acid, with the chemical formula $C_{2}H_{2}O_{4}$. The oxalic acid molecule consists of two carbon atoms, two hydrogen atoms, and four oxygen atoms. In the structural formula, the two carbon atoms are connected by single bonds, and each carbon atom is attached to a hydroxyl group ($OH$) and a carbonyl group ($C=O$), which are the typical structural features of oxalic acid.
Figure 2: A GPT-4o Prompt for Knowledge-Rich Captions. The portion highlighted in red within the prompt emphasizes the instructions for knowledge injection. This prompt utilizes a three-step procedure, which is intended to help ensure the inclusion of a rich body of world knowledge in the generated caption.
By constructing SFT data with captions rich in descriptive details, we enable the model, through fine-tuning, to learn to produce detailed and specific descriptions. Therefore, we designed a detailed prompt and used GPT-4o to generate the initial training data. While GPT-4o understands the prompt requirements, the output data still contain a significant number of hallucinations. To address this, we conducted a manual review of the generated captions. By integrating these data processing steps, we constructed the knowledge-enhanced supervised fine-tuning data suitable for training models capable of generating rich image captions, such as the "arctolamia gestro" illustrated in Figure 1.
@@ 3.3 Training Pipeline
As mentioned above, our comprehensive training pipeline consists of two sequential stages: knowledge-enriching SFT followed by continuous DPO. In this section, we introduce these stages specifically. The general pipeline of our work is shown in Figure 3. We adopt Qwen2-VL-7B (Wang et al., 2024) as the base model and use the LLaMA Factory (Zheng et al., 2024) platform to carry out the overall training process.
Figure 3: The illustration of our recaptioning pipeline. The pipeline comprises the following key stages: Initial Data Generation with GPT-4o, Manual Review, Knowledge-Enriching SFT, and Continuous DPO.
Knowledge-enriching SFT. We sampled raw image-text pairs from open-source multimodal datasets including CC3M (Sharma et al., 2018), CC12M (Changpinyo et al., 2021), DataComp (Gadre et al., 2023), Wukong (Gu et al., 2022), and Wikipedia (Srinivasan et al., 2021). Using the prompt in Figure 2, we employed GPT-4o to generate the initial recaption data. Following a meticulous manual review process, $43,408$ data entries were retained. Subsequently, we performed LoRA fine-tuning on the Qwen2-VL-7B model. Specifically, training was conducted with a learning rate of $51.0e-$ 5, a global batch size of $128$, and for $10$ epochs. This process yielded our Recaption-SFT model.
Continuous DPO. As outlined in Section 3.1, our methodology begins with training the initial DPO model. We first randomly sample $300,000$ image-text pairs and then employ the SFT model for inference. To generate diverse candidates, we used the following sampling hyperparameters: TOPP=$1.0$, TOPK=$20$, and TEMPERATURE=$1.0$. Each image-text pair underwent eight parallel inference iterations, yielding eight divergent candidates. We then constructed preference pairs (chosen/rejected pairs) by selecting the best and worst outputs from each set using an internal critic model.   [^1] This expert system evaluates caption quality by dissecting vision details in captions with an fine-tuned LLM (72B-sized) and assessing hallucination levels for each vision detail with a fine-tuned VLM (72B-sized).. To eliminate length-induced training bias, we implement length-balanced sampling that retains $218k$ high-quality pairs. This process ensures alignment in the length distribution between preferred and dispreferred completions. Subsequently, this curated data set was used for DPO training on the SFT model, yielding the initial DPO model (used as a reference model). Utilizing the trained initial DPO model, we sampled additional $200k$ image-text pairs from the data sources. These pairs were used to generate new preference data employing the model itself, the same sampling hyperparameters, and the critic model. Applying the length-balancing strategy to filter these generated pairs, we retained $139k$ preferred/dispreferred pairs for the subsequent CDPO training iteration. All hyperparameters at different stages are listed in Table 2.
@@ 4 Experiments
This section covers the quality assessment of our recaption data. We first outline the criteria used for this assessment, and then evaluate the data across two dimensions: its intrinsic quality and its impact on large-scale vision language model pre-training.
@@
Figure 4: Performance of different DPO strategies. The horizontal axis represents the training data scale, while the vertical axis depicts the proportion of hallucination-free captions on the validation set (in percentage). Initially, the CDPO strategy utilizes the same training data as DPO. Upon reaching a training scale of $218k$ data points, it incorporates an additional $139k$ new data instances.
@@ 4.1 Evaluation Metrics
The quality of recaption data is assessed along two key dimensions. One dimension concerns the presence of harmful information, particularly hallucinations, which must be minimized. The other is content richness, whose value is validated by its effectiveness in model training. Consequently, our primary quality control efforts focus on mitigating hallucinations. The richness of the data, conversely, is evaluated through the training performance of large-scale vision language models, as detailed in Section 4.3. Adopting the hallucination assessment method from CIEM (Hu et al., 2023), we utilize the large language model, GPT-4o, to parse each caption into its constituent visual details. GPT-4o then assesses each extracted detail for the presence of hallucinations. Following this detail-level judgment, we count the total number of hallucinatory details within a caption. Based on this count, we define two key metrics: the non-hallucination rate as the proportion of captions containing no hallucinatory details and the low-hallucination rate as the proportion of captions containing at most two hallucinatory details.
@@ 4.2 Performance of Recaption Model
To evaluate our model, we apply it to the identical image-text pairs used by other recaption methods, such as Capfusion (Yu et al., 2024) and Recap-DataComp-1B (Li et al., 2024b). The evaluation results are presented in Table 1. When generating data for these same input pairs, our model produces synthetic captions with significantly more details than those generated by Capfusion and substantially fewer hallucinations than those from Recap-DataComp-1B. Specifically, when evaluated on the same data, our model demonstrates substantial improvements over the baselines. Compared to Capfusion, our generated captions exhibit a $267\%$ increase in detail count, are $3.5$ times longer on average, and show a $8.6\%$ percentage point decrease in the rate of hallucinatory details. Relative to Recap-DataComp-1B, our model achieves a $43.6\%$ percentage point increase in the non-hallucination rate. It is important to note that while the brevity of Capfusion’s captions leads to a relatively high non-hallucination rate, the limited detail content restricts their overall value.
We then report the evaluation results of our recaption model on another open source dataset, Wukong (Gu et al., 2022). As illustrated in Figure 4, the base model (Qwen2-VL-7B) exhibits a significant hallucination problem in this dataset. Specifically, its non-hallucination rate is $48.33\%$ and its low-hallucination rate is $87.87\%$, figures that indicate a substantial level of hallucinatory content. Following supervised fine-tuning, the low-hallucination rate is significantly increased. Subsequently, the application of DPO leads to substantial improvements in both the non-hallucination rate and the low-hallucination rate, increasing them $29.53\%$ and $10.21\%$, respectively. Furthermore, we observe that DPO exhibits a notable scaling law during the initial training stage for the recaption task; however, performance plateaus once the data size reaches a certain scale. Our CDPO algorithm provides further enhancement to both metrics, increasing the non-hallucination rate by $3.4\%$ and the low-hallucination rate by $0.5\%$.
@@
Table 3: Experimental results on 15 vision-language tasks. Avg. denotes the average score across all evaluated vision-language tasks. The Non-Hallucination Rate is a hallucination assessment metric defined in Section 4.1. For evaluation, we randomly sampled 1000 instances from the DataCamp dataset to quantify the non-hallucination rate. In the table, bold text indicates the optimal performance for a given metric under the $20M$ data setting, while underlined text signifies further enhancement achieved when utilizing the entire dataset. The $\ast$ symbol indicates that a light, unified supervised fine-tuning (SFT) (Wang et al., 2023; Li et al., 2024a) was applied after pre-training. It should be noted that our primary objective is to investigate the impact of different caption datasets on VLM model training. Consequently, these results represent comparative experiments conducted at a fixed data scale ($20M$). They were performed without employing advanced fine-tuning or reinforcement learning techniques often used to boost performance towards state-of-the-art, and thus are not directly comparable to SOTA models.
@@ 4.3 Performance of Recaption Data
Vision-Language Model Pre-Training. We first validate the effectiveness of our dataset for the pre-training of vision language models. This aims to demonstrate that, compared to existing alt-text and other recaption datasets, Hunyuan-Recap100M can facilitate more effective training of vision language models, leading to superior performance on both general open-source multimodal tasks and our internal knowledge-related test sets. For the VLM framework, we adopt the LLaVA (Liu et al., 2023) architecture, comprising SigLIP (Zhai et al., 2023a) as visual encoder and Hunyuan-7B (Hunyuan, 2025) as the base large language model. To ensure a fair comparison of the influence of caption datasets, we sample $20M$ data points from each of the alt-text and recaption datasets. Note that the experiments that use the alt-text data utilize the same image set as our Hunyuan-Recap100M dataset. In contrast, the Recap-Datacamp-1B data uses a randomly sampled subset, implying a different image set distribution compared to the alt-text and Hunyuan-Recap100M data. [^2] Hunyuan-Recap100M dataset is compiled from multiple open-source datasets, resulting in a distribution that is not exactly the same as that of DataCamp. In Table 3, "Qwen2-VL-7B(20M)" refers to the model trained on the $20M$ recaption data generated by the unoptimized Qwen2-VL-7B model, which contains more hallucinations. On 15 open-source multimodal tasks, the VLM model pre-trained with Hunyuan-Recap100M data significantly outperforms models pre-trained with alt-text or other synthetic captions. This superiority is particularly evident in hallucination-related tasks, where training with Hunyuan-Recap100M data achieves a substantial lead over training with other counterparts. Additionally, we monitored the loss function to evaluate the model convergence behavior in the appendix A. When subjected to identical supervised fine-tuning datasets, models pre-trained with Hunyuan-Recap100M demonstrated markedly improved convergence properties.
Furthermore, to quantify the enhancement in the visual perception capabilities of VLM attributable to Hunyuan-Recap100M, we manually collected $1,197$ images with diverse visual characteristics across 20 domains, including animals, plants, clothing, and landmarks, among others. These images are accompanied by manually annotated labels for evaluation. We assessed the VLM’s perceptive ability on this dataset, and the results are presented in Table 4. When trained with a $20M$ subset of Hunyuan-Recap100M data, the VLM model significantly outperforms other counterparts in 14 of the 20 categories, demonstrating an average improvement of at least $3.8\%$ compared to models trained on other $20M$ datasets. Upon utilizing the full Hunyuan-Recap100M dataset for training, the model shows significant improvements in 18 of the 20 perception tasks, with an overall average improvement reaching $9\%$.
@@
Alt-text(20M) 54.7 36.83 22.67 24.13 29.93 37.07 25.30 50.57 29.47 26.73 17.10 41.80 60.57 37.77 24.73 67.77 41.30 50.30 28.17 31.00 36.33
Recap-DataComp-1B(20M) 32.57 26.60 20.87 21.53 34.47 30.07 24.83 46.57 19.47 10.20 14.73 25.50 54.57 28.57 26.93 60.67 24.63 36.13 37.60 21.93 29.67
Qwen2-VL-7B(20M) 45.83 35.60 34.47 31.90 49.77 45.73 34.80 66.10 36.30 23.83 21.40 40.60 54.30 32.17 29.43 69.83 39.73 43.87 28.90 36.53 39.97
Hunyuan-Recap100M(20M) 42.80 45.30 27.10 32.03 43.90 54.43 34.80 68.23 37.83 28.73 28.13 52.70 61.83 44.10 38.27 74.30 34.00 45.60 45.13 35.23 43.80
Hunyuan-Recap100M 53.43 52.10 39.03 46.07 58.70 55.40 52.53 74.57 44.57 40.47 33.13 67.10 75.27 62.60 54.07 83.87 41.50 50.87 41.63 43.53 53.67
Table 4: Evaluation results on the identification of 20 common visual object categories. Text in bold indicates the optimal performance for a given category under the $20M$ data setting, while underlined text signifies further performance enhancement achieved when utilizing the entire dataset. All reported scores were generated by GPT-4o on May 6, 2025. To mitigate scoring fluctuations, we conducted three independent scoring iterations and used the mean value as the final score.
Text-to-Image Model Training We also evaluate the effectiveness of Hunyuan-Recap100M on the text-to-image generation task. By fine-tuning the Hunyuan-DiT (Li et al., 2024d), we aim to demonstrate that our data can improve visual generation capabilities. As evidenced in Table 5, models fine-tuned from the publicly released Hunyuan-DiT using our recaption data significantly achieve a lower FID and a higher CLIP score on the MSCOCO dataset and an internal perceptual test set. These findings indicate that our data offer considerable assistance in enhancing the model’s understanding of visual concepts and enabling text-to-image models to achieve better performance in terms of authenticity. Realistic cases are presented in Section B.
@@
Table 5: Text-to-Image Evaluation of the Hunyuan-DiT model on an internal perceptual test set and MSCOCO-30K, comparing performance with and without fine-tuning using our recaption data. It should be noted that, due to computational resource limitations, we only sampled $2M$ data points from the Hunyuan-Recap100M dataset for LoRA fine-tuning.